The number ways to select 4 dishes out of 6 different dishes is :
(1) 30 (2) 24
(3) 4 (4) 15

Answer by Maths68(1474)

Use formula = nCr=n!/(n-r)!r!
The number ways to select 4 dishes out of 6 different dishes =6C4
=6!/(6-4)! 4!
=6!/(2)! 4!
=6*5*4*3*2*1/(2*1) 4*3*2*1
=6*5/2
=3*5
=15